# Learning Adaptive and Temporally Causal Video Tokenization in a 1D Latent Space

### v1

@@ 4.3 Ablation Studies
Adaptive training and inference paradigms. AdapTok leverages two adaptive mechanisms: a block-mask sampler for training and an adaptive scorer for inference. The block-mask sampler enables the model to learn from variable token lengths by randomly sampling token counts per block, while the adaptive scorer allocates tokens based on reconstruction quality scores during inference. As shown in Tab. 3a, leveraging both mechanisms results in the best performance, achieving the lowest rFVD of 36.36. The performance gain becomes more pronounced with lower token budgets in inference, highlighting the importance of our adaptive mechanisms for efficient video representation.
Table 3: Ablation studies. (a) Ablation on adaptive training and inference paradigms. (b) Evaluation on different token allocation strategies. (c) Comparison on token allocation strategies.
Tokens Mask Sampler Scorer rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
(a) Adaptive paradigms.
@@
(b) Token allocation strategies.
@@
(c) Scoring metrics.
Token allocation strategies. We explore several token allocation strategies: 1) Fixed token allocation, which assigns a fixed token count to all blocks; 2) Score-threshold binary search (BiThr), which binary searches for a global score threshold and selects the minimal token count per block with predicted score below it; 3) Delta-score binary search (BiDelta), which selects the minimal token count where the score improvement between consecutive lengths drops below a binary searched delta; and 4) Integer Linear Programming (ILP), which jointly optimizes token allocation across a mini-batch under an average token budget. As reported in Tab. 3b, the ILP strategy outperforms the other methods across all metrics. For a comparison under varying token lengths, refer to Fig. 4.
Figure 4: Comparison of token allocation strategies. ILP achieves the best performance across four metrics: (a) FVD, (b) PSNR, (c) MSE, and (d) LPIPS.
Token allocation scoring metrics. We also investigate various scoring metrics for the token allocation, including per-sample SSIM, PSNR, MSE, and perceptual loss. As shown in Tab. 3c and Fig. 5, each scoring metric performs best on its corresponding evaluation metric, e.g., PSNR score yields the highest PSNR (Fig. 5b), MSE score leads to the lowest MSE (Fig. 5c). Notably, using perceptual loss as the scoring metric also leads to superior performance on other evaluation metrics, such as rFVD (Fig. 5a), indicating its stronger correlation with overall perceptual quality.
Figure 5: Comparison of scoring metrics. Perceptual loss achieves better overall performance, especially on FVD and LPIPS.
@@ 4.4 Visualizations
Figure 6: Reconstruction comparison with ElasticTok yan2024elastictok under varying token budgets. AdapTok reconstructs videos in a coarse-to-fine manner, where early tokens capture global structure and later ones refine local details, while ElasticTok suffers from blocky artifacts.
Head tokens encode global information. The tail token drop strategy encourages the head tokens to capture more global information. As shown in Fig.7, attention maps reveal that early tokens attend broadly across the entire spatial field, indicating their role in modeling global semantics. In contrast, mid-to-late tokens focus on localized regions, refining finer details.
Figure 7: Attention maps for latent tokens. Each map shows the attention distribution of a token over spatial patches, with tokens ordered from top-left to bottom-right. Early tokens capture global context, while later tokens focus on local details.
@@ 5 Conclusion
Limitations and Future work. In this work, we focus on the design of a sample-wise, content-aware, and temporally adaptive tokenization framework. Our current implementation adopts a discrete tokenizer based on VQ-VAEs van2017vqvae ; esser2021vqgan , and future work may explore continuous alternatives to validate the generality of the proposed framework. In addition, our model is trained on less than 0.5M publicly available videos due to limited computational resources. Future efforts will focus on scaling up model capacity and dataset diversity to improve generalization and applicability across broader domains.
@@ Appendix A Implementation Details
The detailed training hyper-parameter settings for the AdapTok, Scorer, AdapTok-AR (class-conditional generation on UCF-101) and AdapTok-FP (frame prediciton on Kinetics-600) are reported in Tab. 4.
@@
Appendix B Adaptive Inference
@@
1 Inputs: Video $x$, Block idx $i$;
@@ 13            $s_{max}=s_{mid}$;
Appendix C Visualizations
C.1 Video Reconstruction
@@ C.2 Video Generation
C.3 Token Contribution Analysis

### v2

@@ 4.3 Ablation Studies
Adaptive training and inference paradigms. AdapTok leverages two adaptive mechanisms: a block-mask sampler for training and an adaptive scorer for inference. The block-mask sampler enables the model to learn from variable token lengths by randomly sampling token counts per block, while the adaptive scorer allocates tokens based on reconstruction quality scores during inference. As shown in Tab. 4, leveraging both mechanisms results in the best performance, achieving the lowest rFVD of 36.36. The performance gain becomes more pronounced with lower token budgets in inference, highlighting the importance of our adaptive mechanisms for efficient video representation.
Table 4: Ablation on adaptive training and inference.
Tokens Sampler Scorer rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
[…]
Method Time (ms/video)
ElasticTok 571.7
AdapTok 50.9
Table 6: Ablation on token allocation strategies.
@@
Table 7: Comparison on scoring metrics.
@@
Token allocation strategies. We explore several token allocation strategies: 1) Fixed token allocation, which assigns a fixed token count to all blocks; 2) Score-threshold binary search (BiThr), which binary searches for a global score threshold and selects the minimal token count per block with predicted score below it; 3) Delta-score binary search (BiDelta), which selects the minimal token count where the score improvement between consecutive lengths drops below a binary searched delta; and 4) Integer Linear Programming (ILP), which jointly optimizes token allocation across a mini-batch under an average token budget. As reported in Tab. 6, the ILP strategy outperforms the other methods across all metrics. Fig. 4 further compares the different token allocation strategies across varying token lengths, demonstrating that the ILP-based allocation strategy consistently outperforms the alternatives.
Figure 4: Comparison of token allocation strategies. ILP achieves the best performance across all metrics: (a) FVD, (b) PSNR, and (c) LPIPS.
Token allocation scoring metrics. We also investigate various scoring metrics for the token allocation, including per-sample SSIM, PSNR, MSE, and perceptual loss. As shown in Tab. 7, each scoring metric performs best on its corresponding evaluation metric, e.g., PSNR score yields the highest PSNR. Notably, using perceptual loss as the scoring metric also leads to superior performance on other evaluation metrics, such as rFVD, indicating its stronger correlation with overall perceptual quality. For a comparison under varying token lengths, please refer to the Appendix C.
[…]
Figure 5: Attention maps for latent tokens. Each map shows the attention distribution of a token over spatial patches, with tokens ordered from top-left to bottom-right.
@@ 4.4 Visualizations
Head tokens encode global information. The tail token drop strategy encourages the head tokens to capture more global information. As shown in Fig.5, early tokens capture global context, while later tokens focus on local details. Please refer to the Appendix D for more visualization results.
@@ 5 Conclusion
Limitations and Future work. In this work, we focus on the design of a sample-wise, content-aware, and temporally adaptive tokenization framework. Our current implementation adopts a discrete tokenizer based on VQ-VAEs (Van Den Oord et al., 2017; Esser et al., 2021), and future work may explore continuous alternatives to validate the generality of the proposed framework. In addition, our model is trained on less than 0.5M publicly available videos due to limited computational resources. Future efforts will focus on scaling up model capacity and dataset diversity to improve generalization and applicability across broader domains.
Ethics Statement
This work uses only publicly available datasets that have been rigorously filtered to mitigate potential biases and ethical concerns.
Reproducibility Statement
The implementation details of our method are fully described in Sec. 4.1 and Appendix. A, including both hyper-parameter settings and training costs. Additionally, the pseudo-code for IPAL and details of other token allocation strategies are provided in Algorithm 1 and Appendix. B. The source code and checkpoints will be released to enable reproduction of the main results presented in this paper.
@@ Appendix A Implementation Details
The detailed training hyper-parameter settings for the AdapTok, Scorer, AdapTok-AR (class-conditional generation on UCF-101) and AdapTok-FP (frame prediciton on Kinetics-600) are reported in Table 8. The architectural configurations of the different AdapTok variants are listed in Table 9.
@@
Table 9: Model configurations of AdapTok variants.
Model Hidden Size Depth Heads Parameters
AdapTok-S 512 6 8 59M
AdapTok-L 768 12 12 259M
AdapTok-XL 1024 24 16 913M
Appendix B More Technical Details
B.1 Quantization method
Following Wang et al. (2024a), a stochastic vector quantization (SVQ) is adopted to the quantizer $\mathcal{Q}$. Similar to VQ, a codebook $C\in\mathbb{R}^{c\times d^{\prime}}$ containing $c$ codes is maintained. Given a visual feature $z$, the cosine similarity with all code vectors in $C$ is computed, followed by a softmax to obtain a probability distribution. An index $x_{\mathrm{Ind}}$ is then sampled from this distribution via a categorical distribution:
Given the sampled index, the quantized feature $z_{q}$ is retrieved from the codebook, i.e., $z_{q}=C_{x_{\mathrm{Ind}}}$. To enable differentiable training, the straight-through estimator Bengio et al. (2013) is applied.
B.2 Adaptive Inference
@@
1Inputs: Video $x$, Block idx $i$;
@@ 13    $s_{max}=s_{mid}$;
Appendix C Additional Experiments
Ablation on token allocation scoring metrics.
Fig. 6 presents detailed comparisons of token allocation under varying token lengths for different scoring metrics. Each metric achieves the best performance on its corresponding evaluation metric, while perceptual loss consistently yields strong results across multiple metrics, demonstrating its effectiveness for guiding adaptive token allocation.
Figure 6: Comparison of scoring metrics. Using perceptual loss as the scoring metric achieves better overall performance, especially on FVD and LPIPS.
Further analysis on computational costs.
Although IPAL is based on ILP, its actual runtime overhead is not significant. Table 10-12 shows the runtime scalability of IPAL w.r.t. mini-batch size, the number of blocks and average token numbers. Specifically, across batch sizes ranging from 8 to 1024, FVD remains stable between 36 and 37, while IPAL accounts for only about 15% of total inference time. The allocation time also remains stable with varying average token numbers, while it increases moderately as the number of blocks grows. These results suggest that our method is efficient, scalable, and practical for real-world deployment scenarios.
Table 10: Ablation on mini-batch size with an average token of 512.
Batch size FVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Time (ms/video) IPAL Time
Total IPAL Proportion (%)
8 60.39 24.01 0.146 66.3 28.0 42.2
16 60.14 24.04 0.145 53.7 16.7 31.2
64 59.96 24.06 0.144 45.2 8.6 19.1
128 60.61 24.07 0.144 44.1 7.0 15.8
256 61.48 24.08 0.144 44.0 6.7 15.2
512 61.44 24.08 0.144 43.0 6.3 14.8
1024 61.37 24.09 0.144 44.1 6.8 15.5
Table 11: Ablation on mini-batch size with an average token of 1024.
Batch size FVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Time (ms/video) IPAL Time
Total IPAL Proportion (%)
8 36.55 25.68 0.115 69.0 28.1 40.7
16 36.80 25.71 0.114 56.8 16.9 29.8
64 36.36 25.72 0.114 48.1 8.5 17.8
128 36.52 25.73 0.114 46.7 7.2 15.4
256 37.02 25.74 0.114 47.4 6.9 14.6
512 36.93 25.74 0.113 46.9 7.0 14.8
1024 37.01 25.75 0.113 47.3 7.5 15.8
Table 12: Runtime ablation of IPAL on block and token counts. (a) Ablation on block numbers. (b) Ablation on average token counts.
Blocks Tokens IPAL Time (ms/video)
1 1024 7.9
2 1024 7.1
4 1024 8.0
8 1024 10.6
(a) Number of blocks.
Blocks Tokens IPAL Time (ms/video)
4 32 7.8
4 128 8.5
4 512 8.3
4 1024 8.0
(b) Average tokens counts.
Appendix D More Visualizations
D.1 Video Reconstruction
Figure 7: Reconstruction comparison with ElasticTok Yan et al. (2024) under varying token budgets. AdapTok reconstructs videos in a coarse-to-fine manner, where early tokens capture global structure and later ones refine local details, while ElasticTok suffers from blocky artifacts.
@@ D.2 Video Generation
D.3 Token Contribution Analysis
@@ D.4 Scorer Predictions
Appendix E The Use of Large Language Models (LLMs)
We used LLMs solely to assist with grammar correction of the submission. All research ideas, experiments, and analyses presented in this work were conceived, conducted, and verified by the authors.